-5-(3x+9)=-5 (you distribute the negative sign to all the numbers after it)
-5-3x-9=-5 (after, you combine like terms)
-14-3x=-5 (now, you add 14 and 5, but dnt forget to keep both negative signs)
-3x=-19 (divide both numbers by -3)
x=6 1/3 (now this would be your answer)

Answer by Josellyn430(1)

-5-(3x+9)=-5 (you distribute the negative sign to all the numbers after it)
-5-3x-9=-5 (after, you combine like terms)
-14-3x=-5 (now, you add 14 and 5, but dnt forget to keep both negative signs)
-3x=-19 (divide both numbers by -3)
x=6 1/3 (now this would be your answer)